# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

### v1

@@ Abstract
Scene graph generation (SGG) involves analyzing images to extract meaningful information about objects and their relationships. Given the dynamic nature of the visual world, it becomes crucial for AI systems to detect new objects and establish their new relationships with existing objects. To address the lack of continual learning methodologies in SGG, we introduce the comprehensive Continual ScenE Graph Generation (CSEGG) dataset along with 3 learning scenarios and 8 evaluation metrics. Our research investigates the continual learning performances of existing SGG methods on the retention of previous object entities and relationships as they learn new ones. Moreover, we also explore how continual object detection enhances generalization in classifying known relationships on unknown objects. We conduct extensive experiments benchmarking and analyzing the classical two-stage SGG methods and the most recent transformer-based SGG methods in continual learning settings, and gain valuable insights into the CSEGG problem. We invite the research community to explore this emerging field of study. All data and source code are publicly available at here.
@@ 1 Introduction
Scene graph generation (SGG) aims to extract object entities and their relationships in a scene. The resulting scene graph, carrying semantic scene structures, can be used for a variety of downstream tasks such as object detection(Szegedy et al., 2013), image captioning (Hassan et al., 2023; Aditya et al., 2015) , and visual question answering (Ghosh et al., 2019). Despite the notable advancements in SGG, current works have largely overlooked the critical aspect of continual learning. In the dynamic visual world, new objects and relationships are introduced incrementally, posing challenges for SGG models to adapt and generalize without forgetting previously acquired knowledge. This problem of Continual ScenE Graph Generation (CSEGG) holds great potential for various applications, such as real-time robotic navigation in dynamic environments and adaptive augmented reality experiences.
(a) Scene Graph Generation (SGG)
(b) Continual ScenE Graph Generation (CSEGG)
@@
Due to a scarcity of research specifically addressing the challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable CSEGG. While the field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification (Mai et al., 2021) , object detection (Wang et al., 2021a) , and visual question answering (Lei et al., 2022), these endeavors have largely neglected the distinctive complexities associated with CSEGG. Factors such as the long-tailed distribution of objects and relationships from each task in CSEGG and the intricate interplay in forgetting and generalization between incremental object detection and relationship classification remain unaddressed by previous works.
In this study, we establish a methodology for investigating CSEGG. Building upon existing SGG datasets (Krishna et al., 2017; Kuznetsova et al., 2020), we contribute a comprehensive CSEGG dataset containing 11 tasks, 108,249 images, 150 object classes, and 50 relationships, along with 3 learning protocols, 8 metrics, and a benchmark of continual learning baselines. Our research focuses on analyzing the extent to which existing scene graph generation methods, when combined with common continual learning techniques, experience forgetting as they learn new object entities and relationships. Additionally, we assess the generalization ability of CSEGG models in classifying known relationships between unknown objects. We expand our discussions on the problem motivation in Sec.A.1.4 and highlight the existence of a significant performance gap in this field. We believe that our work provides a valuable testbed for the research community to explore and advance this emerging area of study.
Main Contributions
1. Building upon existing SGG datasets, we introduced a large CSEGG dataset containing 108,249 images, 150 object classes, and 50 relationships.
2. We designed a systematic methodology for conducting CSEGG experiments and evaluating CSEGG models over a sequence of SGG tasks across 3 learning protocols. This helps the community expand the studies of CSEGG and benchmark future AI models.
3. In the experiments, we investigated the unique long-tailed challenge in CSEGG and observed that existing SGG models with continual learning methods do not perform well.
4. We assessed the ability of CSEGG models to generalize and classify known relationships on unfamiliar objects. Our findings revealed that continual learning models exhibiting less forgetting demonstrated better generalization capabilities in this regard.
@@ 2 Related Works
Visual Phrase (Sadeghi & Farhadi, 2011) stands as one of the earliest datasets in the field of visual phrase recognition and detection. Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) (Johnson et al., 2015; Lu et al., 2016; Krishna et al., 2017; Kuznetsova et al., 2020; Liang et al., 2019; Zareian et al., 2020; Yang et al., 2019; Xu et al., 2017; Zhang et al., 2017; Dai et al., 2017; Li et al., 2017b; Zhang et al., 2019a). Among these, the Visual Genome dataset (Krishna et al., 2017) has played a pioneering role by providing comprehensive annotations of objects, attributes, and relationships in images. Another notable dataset is the Visual Relationship Detection dataset (Lu et al., 2016), which specifically focuses on detecting and classifying object relationships. More recently, the Open Image V6 dataset (Kuznetsova et al., 2020) has contributed to scene graph generation tasks by offering a vast collection of images accompanied by localized narratives. Despite the significant contributions of these datasets to SGG, none have been explicitly designed for the task of Continual ScenE Graph Generation (CSEGG). To address this gap, we leverage the existing datasets (Krishna et al., 2017) and curate a new one tailored specifically for CSEGG.
@@
Scene Graph Generation Models. SGG models are categorized into two main approaches: top-down and bottom-up. Top-down approaches(Liao et al., 2019; Yu et al., 2017) typically rely on object detection as a precursor to relationship prediction. They involve detecting objects and then explicitly modeling their relationships using techniques such as rule-based reasoning(Lu et al., 2016) or graph convolutional networks (Yang et al., 2018). On the other hand, bottom-up approaches focus on jointly predicting objects and their relationships in an end-to-end manner (Li et al., 2017a; b; Xu et al., 2017). These methods often employ graph neural networks (Li et al., 2021; Zhang et al., 2019b) or message-passing algorithms (Xu et al., 2017) to capture the contextual information and dependencies between objects. Furthermore, recent works have explored the integration of language priors (Plummer et al., 2017; Lu et al., 2016; Wang et al., 2019) and attention mechanisms in transformers(Andrews et al., 2019) to enhance the accuracy and interpretability of scene graph generation. However, none of these works evaluate SGG models in incremental relationship learning and continual adaptation to new object entities. We benchmark SGG models in CSEGG, identify research gaps, and provide valuable insights about the development of future CSEGG models.
@@
Continual Learning Methods. Existing continual learning works can be categorized into several approaches. (1) Regularization-based methods (Chaudhry et al., 2018; Zenke et al., 2017; Aljundi et al., 2018; Benzing, 2022) aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space, such as Elastic Weight Consolidation (EWC)(Kirkpatrick et al., 2017) or Synaptic Intelligence (SI)(Zenke et al., 2017). (2) Replay-based methods (Rolnick et al., 2019; Chaudhry et al., 2019; Riemer et al., 2018; Vitter, 1985; Rebuffi et al., 2017; Castro et al., 2018) utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting. The variants of these methods include generative replays(Shin et al., 2017; Wu et al., 2018; Ye & Bors, 2020; Rao et al., 2019), where synthetic data is generated and replayed. (3) Dynamic architecture-based approaches(Wang et al., 2022a) adapt the model’s architecture dynamically to accommodate new tasks. Techniques like network expansion(Yoon et al., 2017; Hung et al., 2019; Ostapenko et al., 2019) grow network capacity to accommodate additional knowledge without interfering with the existing ones. Despite the extensive investigation of continual learning in domains like image classification (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021) and object detection(Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), there is a notable dearth of research focusing on CSEGG. This work aims to bridge this gap by addressing the distinctive challenges of SGG, including the issue of long-tailed distribution (Desai et al., 2021; Nan et al., 2021; Chiou et al., 2021), within continual learning.
Figure 2: Three learning scenarios are introduced. From left to right, they are S1. relationship (Rel.) incremental learning (Incre.); S2. relationship and object (Rel. $+$ Obj.) Incre.; and S3. relationship generalization (Rel. Gen.) in Object Incre.. In S1 and S2, example triplets in the training (solid line) and test sets (dash line) from each task are presented. The training and test sets from the same task are color-coded. The new objects or relationships in each task are bold and underlined. In S3, one single test set (dashed gray box) is used for benchmarking the relationship generalization ability of object incre. learning models across all the tasks.
3 Continual ScenE Graph Generation (CSEGG) Benchmark
@@
In CSEGG, we consider a sequence of tasks consisting of images and corresponding scene graphs with new objects or new relationships, or both. Let $D_{t}=\{(I_{i},G_{i})\}_{i=1}^{N_{t}}$ represent the dataset at task $t$, where $I_{i}$ denotes the $i$-th image and $G_{i}$ represents the associated scene graph. The scene graph $G_{i}$ comprises a set of object nodes $O_{i}$ and their corresponding relationships $R_{i}$. Each object node $o_{j}$ is defined by its class label $c_{j}$ and its spatial coordinates. Each relationship $r_{k}$ is represented by a triplet $(o_{s},p_{k},o_{o})$, where $o_{s}$ and $o_{o}$ denote the subject and object nodes, respectively, and $p_{k}$ represents the relationship predicate. The goal of CSEGG is to develop models that can incrementally learn new objects, new relationships, or both, without forgetting previously learned knowledge. Next, we introduce three continual learning scenarios.
(a) S1: Rel. Incre.
(b) S2: Rel. $+$ Obj. Incre.
(c) S3 Rel. Gen. In Obj. Incre.
Figure 3: Label distribution in each task in each learning scenario is presented. In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks. The color gradient indicates the frequency of data within a task, with the lighter color denoting the smaller frequency of data in that category. Only the most frequent labels (relationship labels in (a) and object labels in (c)) are provided. See the legend for the total data size per task. In (b) scenario S2 on both objects and relationships, data distributions are presented in the form of small-world networks, where nodes denote object categories and the edges linking object pairs indicate relationships. Thickness in edges implies the diversity of relationships between object pairs. Same color conventions as (a) and (c) are applied. See the legend for triplet sizes. See Fig. S1,S2,S3 in Sec. A.3 for the full statistics of S1-3.
@@ 3.1 Learning Scenarios
We re-organize and divide the data from Visual Genome (Krishna et al., 2017) to cater to three continual learning scenarios below in our CSEGG dataset. The dataset contains 108,249 images, 150 objects, and 50 relationship categories. We follow the standard image splits for training, validation, and test sets (Xu et al., 2017).
Scenario 1 (S1): Relationship Incremental Learning. While existing continual object detection literature focuses on incrementally learning object attributes (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021; Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene. See Sec. A.1.1 for a concrete example application in medical imaging. To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig. 2S1). There are 5 tasks in S1. To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome (Krishna et al., 2017), and form a set of 10 relationship classes for each task. Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale. Different from continual object recognition and detection tasks, it is difficult to find unique images, in which ground truth scene graphs only contain relationships belonging to one specific task. To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec. A.1.5 for the design motivation). The same reasoning applies in S2 and S3. Example relationship classes from each task and their distributions are provided in Fig. 3(a).
Scenario 2 (S2): Scene Incremental Learning. To simulate the real-world scenario when there are demands for detecting new objects and new relationships over time in old and new scenes, we introduce this learning scenario where new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks (Fig. 2S2). See Sec. A.1.2 for two real-world example applications in robot collaborations on construction sites and video surveillance systems. To select the object and relationship classes from the original Visual Genome (Krishna et al., 2017) for S2, we have two design motivations in mind. First, in real-world applications, such as robotic navigation, robots might have already learned common relationships and objects in one environment. Incremental learning only happens on less frequent relationships and objects. (2) Transformer-based AI models typically require large amounts of training data to yield good performances. Training only on a small amount of data from tail classes often leads to close-to-chance performances. Thus, we take the common objects and relationships from the head classes in Visual Genome as one task, while the remaining less frequent objects and relationships from tail classes as the other task. This results in 2 tasks in total with the first task containing 100 object classes and 40 relationship classes. In the subsequent task, the CSEGG models are trained to continuously learn to detect 25 more object classes and 5 more relationship classes. Same as S1, both the object class and relationship class distributions are still long-tailed within a task (Fig. 3(b)).
@@
Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning. We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects. This scenario is designed to investigate whether the CSEGG models can generalize as well as humans. See Sec. A.1.3 for two real-world applications in the deep sea and space explorations for autonomous navigation systems. Specifically, there are 4 tasks in total with each task containing 30 object classes and 35 relationship classes. In each subsequent task, the CSEGG models are trained to continuously learn to detect 30 more object classes and learn to classify the same set of 35 relationships among these objects. The class selection criteria for each task follow the same as S1, where the selections occur uniformly over head, body, and tail classes. Example object classes and their label distributions for each task are provided in Fig. 3(c). Different from S1 and S2, a standalone generalization test set is curated, where the objects are unknown and their classes do not overlap with any object classes in the training set but the relationships among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on the same generalization test sets.
3.2 Scene Graph Generation Backbone
(a) Schematic of SGTR (Li et al., 2022b)
(b) Overview of three continual learning baselines
Figure 4: Introduction to backbone SGG models and continual learning baselines. We use Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) as the backbone SGG model (Sec. 3.2). SGTR consists of four modules indicated by each blue box. Arrows indicate the signal flows among modules. (b) Three continual learning baselines are listed: EWC (Kirkpatrick et al., 2017), Replay (Rolnick et al., 2019), Naive (Sec. 3.3) and PackNet (Mallya & Lazebnik, 2018)(Sec. 3.3). $\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A. The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
We use the state-of-the-art one-stage Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) and the traditional two-stage SGG model (Xu et al., 2017), named as CNN-SGG. CNN-SGG detects objects with Faster-RCNN(Girshick, 2015) backbone and infers their relationships separately via Iterative message passing (IMP)(Xu et al., 2017). For simplicity and consistency, we focus discussions solely on SGTR in the main text. For CNN-SGG, see Sec. A.2.2 for the introduction, training and implementation details over all three learning scenarios. We observed consistent relative CSEGG performance among all continual learning methods across both backbones (Fig. S15 and Sec. A.5).
Different from CNN-SGG, SGTR formulates the task as a bipartite graph construction problem. Briefly, we introduce how SGTR works (Fig. 4a). Given a scene image $I_{i}$, SGTR utilizes a 2D-CNN followed by a transformer-based encoder to extract image features. These features are further incorporated into a transformer-based decoder to predict object and subject nodes $O_{i}$. After that, the object-aware predicate nodes $R_{i}$ are formed based on both image features and object node features. Finally, a bipartite graph $G_{i}$ is constructed to collectively represent the scene with object $o_{i}$ and predicate nodes $r_{k}$, where the correspondence between these nodes is established based on the Hungarian matching algorithm (Kuhn, 1955). All experimental results are based on the average over 3 runs. We adapt the public source codes from (Li et al., 2022b) and (Wang et al., 2021b) for implementations of the continual learning algorithms on SGTR. We use the hyper-parameters provided in the code as the default values. See Sec. A.2.1 for training and implementation details of SGTR. All source code and data are publicly available here.
Techniques for Learning in Long-tailed Distribution. As shown in Fig. 3 and Sec. 3.1, the data in the training set of every task in all learning scenarios is long-tailed. We adopt the three existing techniques to alleviate the problem of imbalanced data distribution during training. LVIS(Gupta et al., 2019) is an image-level over-sampling strategy. The number of image repeats depends on the object classes with minimal frequency over the entire dataset. Bi-level sampling (BLS) (Li et al., 2021) combines image-level oversampling and instance-level undersampling. LVIS is used for image-level oversampling. At instance levels, a drop-out rate is applied. The more frequent an instance belonging to the common classes, the higher the dropout rate. The two-level data re-sampling achieves an effective trade-off between the head and tail classes. Equalized Focal Loss (EFL) (Li et al., 2022a) is an effective loss function, re-balancing the loss contribution of head and tail classes independently according to their imbalance degrees. During the training of SGTR, EFL is enabled all the time, regardless of the learning scenarios or continual learning algorithms.
3.3 Continual Learning Algorithms

### v3

@@ Abstract
Scene graph generation (SGG) analyzes images to extract meaningful information about objects and their relationships. In the dynamic visual world, it is crucial for AI systems to continuously detect new objects and establish their relationships with existing ones. Recently, numerous studies have focused on continual learning within the domains of object detection and image recognition. However, a limited amount of research focuses on a more challenging continual learning problem in SGG. This increased difficulty arises from the intricate interactions and dynamic relationships among objects, and their associated contexts. Thus, in continual learning, SGG models are often required to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding. To systematically explore Continual Scene Graph Generation (CSEGG), we present a comprehensive benchmark comprising three learning regimes: relationship incremental, scene incremental, and relationship generalization. Moreover, we introduce a “Replays via Analysis by Synthesis" method named RAS. This approach leverages the scene graphs, decomposes and re-composes them to represent different scenes, and replays the synthesized scenes based on these compositional scene graphs. The replayed synthesized scenes act as a means to practice and refine proficiency in SGG in known and unknown environments. Our experimental results not only highlight the challenges of directly combining existing continual learning methods with SGG backbones but also demonstrate the effectiveness of our proposed approach, enhancing CSEGG efficiency while simultaneously preserving privacy and memory usage. All data and source code are publicly available at here.
Keywords:
Scene graph generation Continual learning In-context symbolic replay Long tail distribution Compositional scene graphs
@@ 1 Introduction
Scene graph generation (SGG) aims to extract object entities and their relationships in a scene. The resulting scene graph, carrying semantic scene structures, can be used for a variety of downstream tasks such as object detection[55], image captioning [19, 1] , and visual question answering [16]. Despite the notable advancements in SGG, current works have largely overlooked the critical aspect of continual learning. In the dynamic visual world, new objects and relationships are introduced incrementally, posing challenges for SGG models to account for new changes without forgetting previously acquired knowledge. This problem of Continual ScenE Graph Generation (CSEGG) holds great potential for various applications, such as real-time robotic navigation in dynamic environments and adaptive augmented reality experiences.
@@
The field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification [37], object detection [58], and visual question answering [26]. However, these endeavors have largely neglected the distinctive complexities associated with CSEGG. Here, we highlight several unique challenges of CSEGG: (1) In contrast to object detection, SGG involves understanding and capturing the relationships between objects, which can be intricate and diverse. Consequently, in CSEGG, conveying the spatial and semantic relationships between objects demands adaptive reasoning from the dynamic scene. (2) SGG introduces a higher level of combinatorial complexity than object detection and image classification because each detected object pair may have multiple potential spatial and functional relationships. Thus, as new objects are introduced to the scenes, the complexity of relationships among all the objects increases significantly in a non-linear fashion. (3) The long-tailed distribution in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others. Consequently, CSEGG requires the computational models to adapt continually to the evolving long-tailed distributions over different scenes. Due to a scarcity of research specifically addressing these challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable computational models with the ability of CSEGG.
In this study, we re-organize existing SGG datasets [23, 25] to establish a novel and comprehensive CSEGG benchmark with 3 learning protocols as shown in Fig. 2. (S1). Relationship-incremental setting: an SGG agent learns to recognize new relationships among familiar objects within the same scene. (S2). Scene-incremental setting: an SGG agent is deployed in new scenes where it has to jointly learn to detect new objects and classify new relationships. (S3). Relationship generalization setting: an SGG agent generalizes to recognize known relationships among unknown objects, as the agent learns to recognize new objects.
We curate a set of competitive CSEGG baselines by directly combining three major categories of continual learning methods with two SGG backbones and benchmark them in our CSEGG dataset. Their inferior performances show the difficulties of our benchmark tasks, which require the ability to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding. Specifically, the weight-regularization methods fail to estimate the importance of learnable parameters given the complicated model design in SGG backbones. Although image-replay methods retain knowledge from prior tasks through replays, the extensive combinatorial complexity of relationships among objects surpasses the complexity accommodated by a restricted set of replay images with efficient storage. Additionally, none of these baseline methods consider the shifts inherent in long-tailed distributions in dynamic scenes.
To address the CSEGG challenges, we present a method called "Replays via Analysis by Synthesis", abbreviated as RAS. RAS employs scene graphs from previous tasks, breaks them down and re-composes them to generate diverse scene structures. These compositional scene graphs are then used for synthesizing scene images for replays. Due to its nature of symbolic replays, RAS doesn’t require the storage of original images, which often carry excessive and redundant details. This also ensures data privacy preservation and data efficiency. Furthermore, by synthesizing scenes using composable scene graphs, RAS maintains the semantic context and structure of previous scenes and also enhances the diversity of scene generation. To prevent biased predictions stemming from long-tailed distributions, we moderate the distribution of replayed scene graphs by balancing tail and head classes. This ensures a uniform sampling of relationships and objects during replays. Extensive experiments underscore the effectiveness of our approach. Network analysis reveals our crucial design choices that can be beneficial for the future development of CSEGG models.
@@ 2 Related Works
Scene Graph Generation Datasets. Visual Phrase [50] stands as one of the earliest datasets in the field of visual phrase recognition and detection. Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) [21, 36, 23, 25, 32, 69, 65, 63, 71, 11, 31, 72]. Among these, the Visual Genome dataset [23] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images. Despite the significant contributions of these datasets to SGG, none focuses on continual learning in SGG. As the preliminary efforts towards CSEGG, we re-structure the Visual Genome dataset [23] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
@@
Scene Graph Generation (SGG) Models. SGG models are categorized into two main approaches: top-down and bottom-up. Top-down approaches[33, 68] typically rely on object detection as a precursor to relationship prediction. They involve detecting objects and then explicitly modeling their relationships using techniques such as rule-based reasoning[36] or graph convolutional networks [64]. On the other hand, bottom-up approaches focus on jointly predicting objects and their relationships in an end-to-end manner [30, 31, 63]. These methods often employ graph neural networks [29, 73] or message-passing algorithms [63] to capture the contextual information and dependencies between objects. Furthermore, recent works have explored the integration of language priors [42, 36, 60] and attention mechanisms in transformers[3] to enhance the accuracy and interpretability of scene graph generation. However, none of these works evaluate SGG models in the context of continual learning. In our work, we directly combine continual learning methods with SGG backbones and benchmark these competitive baselines in CSEGG. Our results reveal the limitations of these methods and highlight the challenges of our CSEGG learning protocols.
@@
Continual Learning Methods. Existing continual learning works can be categorized into several approaches. (1) Regularization-based methods [22, 9, 70, 2, 4] aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space. (2) Dynamic architecture-based approaches[57, 67, 20, 41] adapt the model’s architecture dynamically to accommodate new tasks without interfering with the existing ones. (3) Replay-based methods [48, 10, 47, 56, 44, 7] utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting. The special variants of these methods include generative replay methods, such as [52, 62, 66, 43], where synthetic data is generated and replayed. Although these generative replay methods, as well as other continual learning methods, have been extensively studied in image classification [8, 61, 37] and object detection[59, 51, 39], few works focus on the challenges in CSEGG, such as adaptive reasoning from the dynamic scenes, the evolving long-tailed distribution across scenes, and the combinatorial complexity involving objects and their multiple relationships. In this work, we introduce a continual learning method, abbreviated as RAS (Replays via Analysis by Synthesis). To address the distinct challenges of CSEGG, RAS involves creating in-context synthetic scene images based on re-composable scene graphs from previous tasks to reinforce continual learning. The components in RAS facilitate memory-efficient training and preserve privacy while maintaining the scene diversity and scene context for SGG in dynamic environments.
Scenarios #Tasks #Objs #Rels
Eval.
metrics
SGG
backbones
CL
baselines
S1 5 All tasks: 150 Each task: 10 F, R, mF, mR, FWT, BWT, R_bbox, Gen_R Transformer based (SGTR) CNN based (TCNN) Joint, Naive, Replays[48], EWC[22], PackNet[38]
S2 2
Task 1: 100
Task 2: 25
Task 1: 40
Task 2: 5
S3 4 Each task: 30 All tasks: 35
Table 1: Overview of three CSEGG learning scenarios. This table summarizes the three learning scenarios in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines. See Sec. 3 for details.
3 Continual ScenE Graph Generation Benchmark
@@
In CSEGG, to cater to the three continual learning scenarios below, we re-organize the Visual Genome [23] dataset and follow its standard image splits for training, validation, and test sets specified in [63]. In each learning scenario, we consider a sequence of $T$ tasks consisting of images and corresponding scene graphs with new objects, or new relationships, or both. Let $D_{t}=\{(I_{i},G_{i})\}_{i=1}^{N_{t}}$ represent the dataset at task $t$, where $I_{i}$ denotes the $i$-th image and $G_{i}$ represents the associated scene graph. The scene graph $G_{i}$ comprises a set of object nodes $O_{i}$ and their corresponding relationships $R_{i}$. Each object node $o_{j}$ is defined by its class label $c_{j}$ and its bounding box locations and sizes $b_{j}$. Each relationship $r_{k}$ is represented by a triplet $(o_{s},p_{k},o_{o})$, where $o_{s}$ and $o_{o}$ denote the subject and object nodes, and $p_{k}$ represents the relationship predicate.
@@ 3.1 Learning Scenarios
Figure 2: Three learning scenarios are introduced. From left to right, they are S1. relationship (Rel.) incremental learning (Incre.); S2. scene incremental learning; and S3. relationship generalization (Rel. Gen.) in Object Incre.. In S1 and S2, example triplet labels in the training (solid line) and test sets (dash line) from each task are presented. The training and test sets from the same task are color-coded. The new objects or relationships in each task are bold and underlined. In S3, one single test set (dashed gray box) is used for benchmarking the relationship generalization ability of object incre. learning models across all the tasks.
Scenario 1 (S1): Relationship Incremental Learning. To uncover contextual information and go beyond studies of object detection and recognition, we introduce this scenario consisting of 5 tasks where 10 new relationship classes are incrementally added in every task (Fig. 2, left; Tab. 1). All object classes and their locations are made known to all CSEGG models over all the tasks. This scenario resembles a human learning scenario where a parent gradually teaches a baby to recognize new relationships among all objects in the same room, focusing on one new relationship at a time during continual learning. This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context. See Sec. 0.A.1.1 for details.
Scenario 2 (S2): Scene Incremental Learning. To simulate the real-world cases when there are demands for detecting new objects and new relationships from old to new scenes, we introduce this scenario where new objects and new relationships are incrementally introduced over tasks (Fig. 2, middle; Tab. 1). There are 2 tasks in total with the first task containing 100 object classes and 40 relationship classes with 25 more object classes and 5 more relationship classes in the second task. This aligns with the real-world use cases where common objects and relationships are learned in the first scene, and incremental learning in the second scene only happens on less frequent relationships and objects. See Sec. 0.A.1.2 for details.
@@
Scenario 3 (S3): Relationship Generalization. Humans have no problem at all recognizing the relationships of unknown objects with other nearby objects. This scenario is designed to investigate the relationship generalization ability of CSEGG models. This capability is essential for real-world implications, such as in robotic navigation where it often encounters unknown objects and requires classifying their relationships. In total, there are four tasks, each introducing an incremental addition of 30 new object classes. All relationship classes are made known to all CSEGG models over all the tasks (Fig. 2, right; Tab. 1). Different from scenarios S1 and S2, a standalone generalization test set is curated, where the objects are unknown but the relationship classes among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on this standalone generalization test set to predict relationships among the unknown objects. See Sec. 0.A.1.3 for details.
Data sampling and distributions. To allocate data for every task of each scenario, we perform the following sampling strategies. In S1 and S3 above, either object or relationship classes are randomly sampled from the Visual Genome dataset and incrementally added to every task. Due to the inherent characteristics of real-world scenes, the long-tailed class distribution is present in S1 and S3. However, in S2, only tail classes are sampled and added in subsequent tasks.
3.2 Competitive CSEGG Baselines
Due to the scarcity of CSEGG works, we contribute a diverse set of competitive CSEGG baselines and implement them on our own. Each CSEGG baseline requires three components: a backbone model for scene graph generation (SGG), a continual learning (CL) method to prevent the SGG model from forgetting, and an optional data sampling technique to deal with imbalanced data at every task for training SGG models. Next, we introduce the 2 SGG backbones, the 5 continual learning methods, and the 5 optional data sampling techniques. See Sec. 0.A.2 for implementation and training details of CSEGG baselines.
SGG Backbones. We use the two state-of-the-art backbones: (1) one-stage Scene graph Generation TRansformer (SGTR) [28] and (2) the traditional Two-stage SGG model (TCNN) [63]. Briefly, SGTR (Fig. S2(a)) uses a transformer-based architecture for image feature extraction and fusion. During training, [28] formulates SGG as a bipartite graph construction and matching problem. In contrast, TCNN detects objects with Faster-RCNN[17] backbone and infers their relationships separately via Iterative message passing [63]. We use public source codes from [28] and [59] with default hyperparameters.